OpenStudy (anonymous):

A committee of 6 people is chosen from 8 students, 10 teachers, and 6 business executives. What is the probability that: a) there is exactly one student on the committee b) there are no business executives on the committee c) there is at least one teacher on the committee

OpenStudy (anonymous):

i'll you try

OpenStudy (anonymous):

one out of 8, the other 5 can be any of the 16 remaining numerator is \(\binom{8}{1}\times \binom{1}{5}\) denominator is \(\binom{24}{6}\)

OpenStudy (anonymous):

*let*

OpenStudy (anonymous):

oops big typo there

OpenStudy (anonymous):

\[\binom{8}{1}\times \binom{16}{5}\]

OpenStudy (anonymous):

OMG THE GREAT SATELLITE. I've seen you so much when I was looking for help, thanks a lot. That's exactly what I had for the answer to a) too. And sourwing, I'm done, this is just to see if I'm correct or not ;)

OpenStudy (anonymous):

Ok so I got 34944 for the answer, the only thing that was messing me up here is the denominator, how do I find it?

OpenStudy (anonymous):

there are 24 people and you're choosing 6

OpenStudy (anonymous):

OH YES \[\left(\begin{matrix}24 \\ 6\end{matrix}\right)\] THANKS AGAIN SOUR

OpenStudy (anonymous):

yes, you can add up the probabilities that have at least 1 teacher. Or you can find the probability that contain NO teacher and subtract from the universal set

OpenStudy (anonymous):

No businessmen would require missing the businessmen in all six picks, which is (18/24)(17/23)(16/22)(15/21)(14/20)(13/19). The problem of at least one teacher can be handled as @sourwing just noted after doing this kind of calculation for the no teacher situation. I am not familiar with the notations used above to explain the probabilities, so my comment may be equivalent to theirs...or even wrong.

OpenStudy (anonymous):

That is also correct douglas, but just so you know, the above notation is for Combinations, where order of the people is not important. \[\left(\begin{matrix}24 \\ 6\end{matrix}\right)\] is 24C6 which means 6 people are being chosen from a total of 24.

OpenStudy (anonymous):

Thank you for explaining it. Regards.